HUBBLE SPACE TELESCOPE DAILY REPORT #2479 PERIOD COVERED: 0000Z (UTC) 10/14/99 - 0000Z (UTC) 10/15/99 Daily Status Report as of 288/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. The observations were executed as scheduled, and no anomalies were reported. 1.2 Completed Three Sets of WF/PC-2 8443 (WF/PC-2 Cycle-8 Supplemental Darks pt1/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no problems were reported. 1.3 Completed STIS/CCD 8409 (STIS CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and to track the evolution of hot columns. The observations were executed as scheduled, and no anomalies were reported. 1.4 Completed STIS/CCD/MA1 8241 (A Snapshot Survey of Interstellar Absorption Lines) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make high resolution Echelle spectrographic observations of the standard ultraviolet star HD92536 as part of a snapshot survey of interstellar absorption lines. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 8148 (Discrete Photometry of Galileo Mission Atmospheric Targets in Jupiter) The WF/PC-2 was used to make photometric observations of Jupiter's atmosphere. This is to investigate the vertical aerosol structure of Jupiter's atmosphere, and its spatial and temporal variability, in conjunction with nearly simultaneous observations by the uniquely well-positioned but data-rate-limited Galileo Orbiter. As described in yesterday's report and HSTAR 7311, there was an approximate 7-minute loss of lock that occurred about one-half hour into this almost nine-hour proposal. The take data flag was down during this period, affecting one observation. A little over three hours later the loss of lock described in 2.1 and HSTAR 7312 occurred, affecting one observation. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.6 Completed WF/PC-2 8444 (WF/PC-2 CYCLE-8 Internal Monitor) The WF/PC-2 was used to make a bias, and an internal flat, observation as part of a Cycle-8 monitoring program to monitor the health of the cameras. A variety of internal exposures are obtained in order to provide a monitor of the integrity of the CCD camera electronics in both bays {gain 7 and gain 15}, a test for quantum efficiency in the CCDs, and a monitor for possible buildup of contaminants on the CCD windows. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed WF/PC-2 8457 (UV Earthflats) The WF/PC-2 was used to make a ultraviolet Earth flat calibration image. This is to improve the quality of pipeline the ultraviolet flatfields. The observations were completed as scheduled, and no problems were reported. 1.8 Completed WF/PC-02 8131 (HST Imaging of Moderate Redshift X-ray Emitting Groups of Galaxies) The WF/PC-2 was used to make observations of the infrared and x-ray emitting galactic cluster Gal-Clus-0723. Most galaxies in the universe, including our own Milky Way, are members of poor groups of galaxies. Recent optical and X- ray observations of nearby groups have revealed considerable new insight into these important systems. The observations were completed as scheduled, and no problems were reported. 1.9 Completed STIS/CCD 8549 (POMS Test Proposal: STIS non-scripted parallel proposal continuation III) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations as part of a non-scripted POMS Test Proposal. The observations were completed as scheduled, and no problems were reported. 1.10 Completed WF/PC-2 8134 (Is the IMF Universal at Low Stellar Masses?) The WF/PC-2 was used to make observations of the cluster NGC-2172. Only the HST can provide the resolution and limiting magnitude required to reach the low-mass end of the IMF at extragalactic distances. The observations were executed as scheduled, and no problems were reported. 1.11 Completed STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. The observations were completed as scheduled, and no problems were reported. 1.12 Completed WF/PC-2 8216 (T Tauri Star Snapshot Survey II. Completing the CTTs Sample) The WF/PC-2 was used to make observations of the T-Tauri class star known as SZ-10 as part of Snapshot Survey II. This database will be of fundamental importance to understanding the nature of planet-forming environments and their frequency in the galaxy. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 5 Successful: 5 As described in HSTAR 7313, the acquisition at 288/012014Z defaulted to fine lock back-up on FGS-3 only when the scan step limit for FGS-1 was exceeded. Three proposals, occurring during the period of the next report, may have been affected: STIS 88342, two iterations of WF/PC-2 8436, and STIS 8408. Scheduled Re-acquisitions: 10 Successful: 10 Per HSTAR 7312, there was a loss of lock at 287/063444Z that lasted until 287/064028Z that occurred during a type 3 profile slew. The proposal detailed in 1.5 was affected. 2.2 FHST Updates: Scheduled: 13 Successful: 13 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared twice. An SI C&DH error was reset at 287/2021Z as directed by ROP NS-5. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH